A man spends half of his salary on household expenses; 1/4th for rent, 1/5th for travel expenses, the man deposits the rest in a bank. If his monthly deposits in the bank amount $50, what is his monthly salary ?

Answer by rfer(16322)

1/4+1/5+1/2=
5/20+4/20+10/20=19/20
50/(1/20)=salary
50/1*20/1=$1000